OpenStudy (rock_mit182):

A, B are matrices nxn; If A is NON-invertible matrix, A times B IS non-invertible as well. Please help me !

OpenStudy (beginnersmind):

Can you use determinants?

OpenStudy (rock_mit182):

oh i see...

OpenStudy (rock_mit182):

if A INVERSE does not exit then |A| =0

OpenStudy (rock_mit182):

So there are two cases when: B inverse exist and B inverse does't exist, in other words: |B| =! 0 OR |B| = 0 in any case : |AB| = |A|*|B|

OpenStudy (rock_mit182):

WHICH alwas have to be zero given the condition of A

OpenStudy (rock_mit182):

right ?

OpenStudy (beginnersmind):

right

OpenStudy (beginnersmind):

last step is to point out det(AB) = 0 -> AB is non-invertible

OpenStudy (rock_mit182):

dude plz help me out with another question

OpenStudy (beginnersmind):

Sure

OpenStudy (rock_mit182):

show that A^2+2A = - I has inverse WHERE: A is nxn matrix; I is the identity matrix

OpenStudy (beginnersmind):

Hm, let me think.

OpenStudy (rock_mit182):

of course

OpenStudy (beginnersmind):

I guess you can do this one with determinants too :)

OpenStudy (rock_mit182):

i guess on this one i only can use propeties of matrix..

OpenStudy (beginnersmind):

Do we need to prove that A has an inverse or that A^2+2A does?

OpenStudy (beginnersmind):

Anyway, I need to go, so I'll leave with a short proof with determinants From A^2 + 2A = -I, we have det(A)*det(A) + 2det(A) + det(I) = 0 substituting u = det(A) u^2 + 2u + 1 = 0, or detA = 1 or detA = -1, so in either case A is invertible. I don't have a non-determinant proof, but here's a hint: If you take the matrix A^2+2A = -I its rank is n. A^2 + 2A = (A +2I)(A), and now you can use some linear algebra that both A and A+2I are rank n.

OpenStudy (beginnersmind):

u^2 + 2u + 1 = 0, or detA = 1 or detA = -1, CORRECTION u = -1, or detA = -1

OpenStudy (rock_mit182):

thanks dude